OpenStudy (anonymous):

e the substitution method to find all solutions of the system of equations. x2 − y = 13 2x2 + 3y = 41

OpenStudy (anonymous):

use*

OpenStudy (anonymous):

are those \[x ^{2}\]

OpenStudy (anonymous):

\[y=-12+x ^{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[2x ^{2} + 3(-13 +2x ^{2)}= 41\]

OpenStudy (anonymous):

\[8x ^{2} - 39 = 41\]

OpenStudy (anonymous):

\[x =\sqrt{80/8}= \sqrt{10}\]

OpenStudy (anonymous):

y= -13 +10 y=-3

OpenStudy (anonymous):

Do you know how to format the answer like this? (x, y) = (_____ ) (smaller x-value) (x, y) = (_____ ) (larger x-value)

OpenStudy (anonymous):

no dude, sorry